# Variational Stochastic Gradient Descent for Deep Neural Networks

### v1

@@ Abstract
Optimizing deep neural networks is one of the main tasks in successful deep learning. Current state-of-the-art optimizers are adaptive gradient-based optimization methods such as Adam. Recently, there has been an increasing interest in formulating gradient-based optimizers in a probabilistic framework for better estimation of gradients and modeling uncertainties. Here, we propose to combine both approaches, resulting in the Variational Stochastic Gradient Descent (VSGD) optimizer. We model gradient updates as a probabilistic model and utilize stochastic variational inference (SVI) to derive an efficient and effective update rule. Further, we show how our VSGD method relates to other adaptive gradient-based optimizers like Adam. Lastly, we carry out experiments on two image classification datasets and four deep neural network architectures, where we show that VSGD outperforms Adam and SGD.
Machine Learning, ICML
@@ 1 Introduction
Nowadays, adaptive gradient-based methods are the leading optimizers in deep learning. The first advancements arrived with the idea of adding momentum to the SGD update rule, resulting in SGD with momentum (SGDM) (Sutskever et al., 2013). The next major breakthrough was Adam (Kingma & Ba, 2015), which uses the first and second momenta to adapt the learning rate and gradients. Currently, Adam is probably the most widely used optimizer in deep learning due to its relative insensitivity to hyperparameter values, and much faster initial progress in training compared to SGD (Sun, 2020).
Recently, there has been an increasing interest in formulating SGD in a probabilistic framework. One example is (Mandt et al., 2017) in which SGD is used as an approximate Bayesian inference algorithm analyzed by relating it to the Ornstein-Uhlbeck process. A different approach utilizes a Bayesian perspective on SGD to model the uncertainty for datastreams (Liu et al., 2024).
@@ 2 Background knowledge
A different line of research focuses on $2^{\text{nd}}$-order optimization methods that require the calculation of Hessian matrices. In the case of deep learning, it is infeasible to calculate Hessian matrices for all layers. Therefore, many researchers focus on approximated $2^{\text{nd}}$-order optimizers. One of the most popular methods is KFAC (Martens & Grosse, 2015). Recently, there have been new developments in this direction (e.g., (Eschenhagen et al., 2023; Lin et al., 2021, 2023)). However, extending K-FAC introduces memory overhead that could be prohibitive, and numerical instabilities might arise in low-precision settings due to the need for matrix inversions or decompositions. Here, we focus on the $1^{\text{st}}$-order optimizer. However, we see great potential for extending our framework to the $2^{\text{nd}}$-order optimization.
There are various attempts to treat SGD as a probabilistic framework. For instance, Mandt et al. (2017) presented an interesting perspective on how SGD can be used for approximate Bayesian posterior inference. Moreover, they showed how SGD relates to a stochastic differential equation (SDE), specifically, the Ornstein-Uhlbeck process. The idea of perceiving SGD as an SDE was further discussed in, e.g., (Yokoi & Sato, 2019) with interesting connections to stochastic gradient Langevin dynamics (Welling & Teh, 2011). Another perspective was on the phrasing SGD as a Bayesian method either for sampling (Mingard et al., 2021), modeling parameter uncertainty for streaming data (Liu et al., 2024) or generalization (Smith & Le, 2017). A different approach uses linear Gaussian models to formulate the generation process of the noisy gradients and infer the hidden true gradient with Kalman filter or Bayesian filter, e.g., (Bittner & Pronzato, 2004; Vuckovic, 2018; Yang, 2021). In this paper, we propose to model dependencies between true gradients and noisy gradients through the Bayesian perspective using a novel framework. Our framework treats both quantities as random variables and allows efficient inference and incorporation of prior knowledge into the optimization process. A similar line of work is the natural gradient-based methods for Bayesian Inference where the optimizer uses natural gradients instead of Euclidean gradients (Khan et al., 2018; Osawa et al., 2019). Our work can be considered orthogonal to this line of work, as they apply Adam to a Bayesian model for learning the variational parameters. VSGD, on the other hand, assumes a Bayesian model for the gradients themselves and may not necessarily be applied to a Bayesian inference problem.
@@ 3.1 Our approach: VSGD
In our considerations, we treat all quantities as random variables. Hence, the true gradient $g_{t}$ is a latent variable, and we observe the noisy gradient $\hat{g}_{t}$. We propose to model $g_{t}$ and $\hat{g}_{t}$ using Gaussian distributions with precision variables $w_{g}$ and $w_{\hat{g}}$, representing the information of systematic and observation noises, respectively. Here, we use ”observation noise” to indicate the sampling noise of the observed gradient and ”systematic noise” to indicate the introduced error of trying to approximate the real, local area of the gradient surface with a simplified model. Both precision variables are treated as Gamma-distributed latent variables. The advantage of treating precision variables as latent variables instead of hyperparameters is that this way, we can incorporate prior knowledge over the systematic and observation noise as well as having uncertainty over them.
We also introduce a control variate $u_{t}$ to serve as a mean for $g_{t}$ and is essential for our approach. We suggest defining it as a function that aggregates information from previously observed noisy gradients $u_{t}=h(\hat{g}_{1:t-1})$. Later (Eq. 8), we will discuss the specific formulation of the function $h(\cdot)$. In the case of $\hat{g}_{t}$, we propose using $g_{t}$ as the mean.
@@
where the Gamma distributions $\Gamma(\cdot)$ are parameterized by shape and rate. $K_{g}$ is a hyperparameter designed to reflect our prior belief about the variance ratio between $\hat{g}_{i}$ and $g_{i}$, i.e., $K_{g}=\mathbb{E}\left[w_{\hat{g}}^{-1}\right]/\mathbb{E}\left[w_{g}^{-1}\right]$. A recommended value for $K_{g}$ is greater than $1$, e.g. $K_{g}=30$, suggesting that the majority of observed variance stems from observation noise $w_{\hat{g}}^{-1}$ rather than systematic noise $w_{g}^{-1}$. $\gamma$ is a shared hyperparameter that indicates the strength of the prior, essentially equating the prior knowledge with $\gamma$ observations. A smaller value for $\gamma$, such as $81e-$ 8, is advised to represent our prior ignorance. The remaining key question is the choice of the control variate $u_{t}$ as a way to carry information. In this paper, we aggregate the information over $t$ by setting the prior mean for the current $g_{t}$ as the posterior mean of the previous ones:
@@
Since the model defined by Eqs. 4-7 requires calculating intractable integral over latent variables, to find a good estimate of $g_{t}$, we aim to approximate the distribution $p(w_{g},w_{\hat{g}},g_{1:T},\hat{g}_{1:T};u_{1:T})$ with a variational distribution $q(w_{g},w_{\hat{g}},g_{1:T};\tau,\phi_{1:T})$, parameterized by the global and local variational parameters $\tau$ and $\{\phi_{t}\}_{t=1}^{T}$ respectively. We achieve this by maximizing the evidence lower-bound (ELBO) defined by $p$ and $q$
@@
Note that $\mu_{t,g}$ represents a dynamic weighted average of the previous estimate $\mu_{t-1,g}$ and the noisy observation $\hat{g}_{t}$. Weights are dynamically determined based on the learned relative precision between Eq. 4 and Eq. 5. $\mu_{t,g}$ then acts as a control variate for the subsequent iteration. Recall that $\cfrac{a_{t,g}}{b_{t,g}}=\mathbb{E}_{q_{t}}[w_{g}]$. Therefore, the variance parameter $\sigma^{2}_{t,g}$ is equal to the inverse sum of the two precisions characterizing the systematic and observation noise. For an alternative kernel smoothing view on Eq. 15, please refer to Appendix A.2.
@@
     Update $\sigma^{2}_{t,g},\mu_{t,g}$triangleright Eq. 15, 16
@@ 3.2 VSGD, Constant VSGD and their connections to other gradient-based deep learning optimizers
If we take $\gamma\rightarrow\infty$, $K_{g}\rightarrow 0$, and $\gamma K_{g}\rightarrow\infty$, then VSGD is approximately equivalent to Normalized-SGD (Murray et al., 2019). We can show it by taking $b_{0,g}=\gamma$, $b_{0,\hat{g}}=K_{g}\gamma$. Then, for $\gamma\rightarrow\infty$ and $K_{g}\rightarrow 0$, Eq. 15 yields the following:
@@
     Update $\sigma^{2}_{t,g},\mu_{t,g}$triangleright Eq. 34, 35
     Update $a_{t,\hat{g}}$, $b_{t,\hat{g}}$triangleright Eq. 36,37
@@
$b_{t,\hat{g}}=$ $(1-\rho_{t})b_{t-1,\hat{g}}+\rho_{t}\Big{[}\gamma$ (37)
$+0.5\left(\sigma^{2}_{t,g}+(\mu_{t,g}-\hat{g}_{t})^{2}\right)$
$+0.5K_{g}\left(\sigma^{2}_{t,g}+(\mu_{t,g}-\mu_{t-1,g})^{2}\right)\Big{]},$
@@ 3.2.1 Constant VSGD and Adam
For clarity, we divide the aforementioned equation into two components: Eq. 46 and Eq. 47. The first component aligns with Eq. 40, as both represent weighted sums of an estimated second momentum and $\hat{g}_{t}^{2}$. The distinction arises in the second component, where Constant VSGD introduces two additional factors into the weighted sum: $\mu_{t-1,g}\hat{g}_{t}$ and $\frac{1}{K_{g}+1}\frac{b_{t-1,\hat{g}}}{a_{t-1}}$. Specifically, $\mu_{t-1,g}\hat{g}_{t}$ applies a penalty on Eq.46 when $\mu_{t-1,g}$ and $\hat{g}_{t}$ have opposing signs. The factor $\frac{1}{K_{g}+1}\frac{b_{t-1,\hat{g}}}{a_{t-1}}$ represents the $\frac{1}{K_{g}+1}$ fraction of observation noise learned from the data, while a higher observation noise implies greater uncertainty in $g_{t}$, leading to a higher expected value $\mathbb{E}\left[g_{t}^{2}\right]$.
In summary, Constant VSGD described in Algorithm 2 can be regarded as a variant of Adam. The first momentum update in both algorithms is equivalent when $K_{g}$ is set to $\frac{\beta_{1}}{1-\beta_{1}}$. However, the second momentum update in Constant VSGD incorporates an additional data-informed component that is dynamically adjusted based on the observation noise learned from the data.
@@ 3.2.2 Constant VSGD and SGD with momentum (SGDM)
$v_{t}$ $=\lambda v_{t-1}+\eta\hat{g}_{t}$ (50)
$=(\lambda+\eta)\left[\frac{\lambda}{\lambda+\eta}v_{t-1}+\frac{\eta}{\lambda+\eta}\hat{g}_{t}\right].$ (51)
That is, the increment $v_{t}$ in SGDM is a weighted average of $v_{t-1}$ and $\hat{g}_{t}$, the counterpart of $v_{t}$ in Constant VSGD is $\mu_{t,g}$ updated by Eq. 34. Therefore, by choosing $K_{g}$ such that $K_{g}=\frac{\gamma}{\eta}$, the update equations 51 and 34 become identical except for a constant scaling factor.
@@ 3.2.3 Constant VSGD and AMSGrad
Similarly to AMSGrad, Constant VSGD incorporates a form of long-term memory regarding past gradient information in its second momentum update equations 46 and 47. However, this is not achieved by computing the maximum of historical values. Notably, the factor $\frac{1}{K_{g}+1}\frac{b_{t-1,\hat{g}}}{a_{t-1}}$ in Eq. 47 represents the expected observation noise, estimated using data up to the $(t-1)$-th iteration. In this expression, while $K_{g}$ and $a_{t-1}$ remain constant for $t\geq 2$, the variable component is $b_{t-1,\hat{g}}$, which is determined by Eq. 37, for clarity we restate it as:
@@
$s=\gamma$ $+0.5\left(\sigma^{2}_{t,g}+(\mu_{t,g}-\hat{g}_{t})^{2}\right)$ (55)
$+0.5K_{g}\left(\sigma^{2}_{t,g}+(\mu_{t,g}-\mu_{t-1,g})^{2}\right).$ (56)

### v2

# algorithmic

@@ Abstract
Optimizing deep neural networks is one of the main tasks in successful deep learning. Current state-of-the-art optimizers are adaptive gradient-based optimization methods such as Adam. Recently, there has been an increasing interest in formulating gradient-based optimizers in a probabilistic framework for better modeling the uncertainty of the gradients. Here, we propose to combine both approaches, resulting in the Variational Stochastic Gradient Descent (VSGD) optimizer. We model gradient updates as a probabilistic model and utilize stochastic variational inference (SVI) to derive an efficient and effective update rule. Further, we show how our VSGD method relates to other adaptive gradient-based optimizers like Adam. Lastly, we carry out experiments on two image classification datasets and four deep neural network architectures, where we show that VSGD outperforms Adam and SGD.
@@ 1 Introduction
Nowadays, adaptive gradient-based methods are the leading optimizers in deep learning. The first advancements arrived with the idea of adding momentum (Polyak, 1964) to the SGD update rule (SGDM) which demostrated impressive performance in training DNNs (Sutskever et al., 2013). The next major breakthrough was Adam (Kingma & Ba, 2015), which uses the first and second momenta to adapt the learning rate and gradients. Currently, Adam is probably the most widely used optimizer in deep learning due to its relative insensitivity to hyperparameter values and much faster initial progress in training compared to SGD (Sun, 2020).
Recently, there has been an increasing interest in formulating SGD in a probabilistic framework. An example is Mandt et al. (2017) in which SGD is used as an approximate Bayesian inference algorithm analyzed by relating it to the Ornstein-Uhlbeck process. A different approach utilizes a Bayesian perspective on SGD to model the uncertainty for datastreams (Liu et al., 2024).
@@ 2 Background knowledge
A different line of research focuses on $2^{\text{nd}}$-order optimization methods that require the calculation of Hessian matrices. In the case of deep learning, it is infeasible to calculate Hessian matrices for all layers. Therefore, many researchers focus on approximated $2^{\text{nd}}$-order optimizers. One of the most popular methods is KFAC (Martens & Grosse, 2015). Recently, there have been new developments in this direction (e.g., (Eschenhagen et al., 2023; Lin et al., 2021; 2023)). However, extending K-FAC introduces memory overhead that could be prohibitive, and numerical instabilities might arise in low-precision settings due to the need for matrix inversions or decompositions. Here, we focus on the $1^{\text{st}}$-order optimizer. However, we see great potential for extending our framework to the $2^{\text{nd}}$-order optimization.
There are various attempts to treat SGD as a probabilistic framework. For instance, Mandt et al. (2017) presented an interesting perspective on how SGD can be used for approximate Bayesian posterior inference. Moreover, they showed how SGD relates to a stochastic differential equation (SDE), specifically, the Ornstein-Uhlbeck process. The idea of perceiving SGD as an SDE was further discussed in, e.g., (Yokoi & Sato, 2019) with interesting connections to stochastic gradient Langevin dynamics (Welling & Teh, 2011). Another perspective was on the phrasing SGD as a Bayesian method either for sampling (Mingard et al., 2021), modeling parameter uncertainty for streaming data (Liu et al., 2024) or generalization (Smith & Le, 2017). A different approach uses linear Gaussian models to formulate the generation process of the noisy gradients and infer the hidden true gradient with the Kalman filter or Bayesian filter, e.g. (Bittner & Pronzato, 2004; Vuckovic, 2018; Yang, 2021). In this paper, we propose to model dependencies between true gradients and noisy gradients through the Bayesian perspective using a novel framework. Our framework treats both quantities as random variables and allows efficient inference and incorporation of prior knowledge into the optimization process. A similar line of work is the natural gradient-based methods for Bayesian Inference, where the optimizer uses natural gradients instead of Euclidean gradients (Khan et al., 2018; Osawa et al., 2019). Our work is different from this line of work, as they apply Adam to learn the variational posteriors of the model. VSGD, on the other hand, learns a variational posterior for the gradients themselves and may not necessarily be applied to estimate the posterior of the model parameters.
@@ 3.1 Our approach: VSGD
In our considerations, we treat all quantities as random variables. Hence, the true gradient $g_{t}$ is a latent variable, and we observe the noisy gradient $\hat{g}_{t}$. We propose to model $g_{t}$ and $\hat{g}_{t}$ using Gaussian distributions with precision variables $w_{g}$ and $w_{\hat{g}}$, representing the information of systematic and observation noises, respectively. Here, we use "observation noise" to indicate the sampling noise of the observed gradient and "systematic noise" to indicate the introduced error of trying to approximate the real, local area of the gradient surface with a simplified model. Both precision variables are treated as Gamma-distributed latent variables. The advantage of treating precision variables as latent variables instead of hyperparameters is that this way, we can incorporate prior knowledge over the systematic and observation noise as well as having uncertainty over them.
We also introduce a control variate $u_{t}$ to serve as the prior mean for $g_{t}$ and is essential for our approach. We define $u_{t}=h(\hat{g}_{1:t-1})$ as a deterministic message passed from previously observed noisy gradients to time $t$. Later (Eq. 7), we will discuss the specific formulation of the function $h(\cdot)$. In the case of $\hat{g}_{t}$, we propose using $g_{t}$ as the mean.
@@
where the Gamma distributions $\Gamma(\cdot)$ are parameterized by shape and rate. $K_{g}$ is a hyperparameter designed to reflect our prior belief about the variance ratio between $\hat{g}_{i}$ and $g_{i}$, i.e., $K_{g}=\mathbb{E}\left[w_{\hat{g}}^{-1}\right]/\mathbb{E}\left[w_{g}^{-1}\right]$. A recommended value for $K_{g}$ is greater than $1$, e.g. $K_{g}=30$, suggesting that the majority of observed variance stems from observation noise $w_{\hat{g}}^{-1}$ rather than systematic noise $w_{g}^{-1}$. $\gamma$ is a shared hyperparameter that indicates the strength of the prior, essentially equating the prior knowledge with $\gamma$ observations. A smaller value for $\gamma$, such as $1$e88-8- 8, is advised to represent our prior ignorance. The remaining key question is the choice of the control variate $u_{t}$ as a way to carry information. In this paper, we aggregate the information over $t$ by setting the prior mean for the current $g_{t}$ as the posterior mean of the previous ones:
@@
Since the model defined by Eqs. 3-6 requires calculating intractable integral over latent variables, to find a good estimate of $g_{t}$, we aim to approximate the distribution $p(w_{g},w_{\hat{g}},g_{1:T},\hat{g}_{1:T};u_{1:T})$ with a variational distribution $q(w_{g},w_{\hat{g}},g_{1:T};\tau,\phi_{1:T})$, parameterized by the global and local variational parameters $\tau$ and $\{\phi_{t}\}_{t=1}^{T}$ respectively. We achieve this by maximizing the evidence lower-bound (ELBO) defined by $p$ and $q$
@@
Note that $\mu_{t,g}$ represents a dynamic weighted average of the previous estimate $\mu_{t-1,g}$ and the noisy observation $\hat{g}_{t}$. Weights are dynamically determined based on the learned relative precision between Eq. 3 and Eq. 4. $\mu_{t,g}$ then acts as a control variate for the subsequent iteration. Recall that $\tfrac{a_{t,g}}{b_{t,g}}=\mathbb{E}_{q_{t}}[w_{g}]$. Therefore, the variance parameter $\sigma^{2}_{t,g}$ is equal to the inverse sum of the two precisions characterizing the systematic and observation noise. For an alternative kernel smoothing view on Eq. 14, please refer to Appendix A.2.
@@
     Update $\sigma^{2}_{t,g},\mu_{t,g}$triangleright Eq. 14, 15
@@ 3.2 VSGD, Constant VSGD and their connections to other gradient-based deep learning optimizers
If we take $\gamma\rightarrow\infty$, $K_{g}\rightarrow 0$, such that $\lim_{\gamma\rightarrow\infty K_{g}\rightarrow 0}\gamma K_{g}=0$, then VSGD is approximately equivalent to Normalized-SGD (Murray et al., 2019). We can show it by taking $b_{0,g}=\gamma$, $b_{0,\hat{g}}=K_{g}\gamma$. Then, for $\gamma\rightarrow\infty$ and $K_{g}\rightarrow 0$, Eq. 14 yields the following:
@@
     Update $\sigma^{2}_{t,g},\mu_{t,g}$triangleright Eq. 33, 34
     Update $a_{t,\hat{g}}$, $b_{t,\hat{g}}$triangleright Eq. 35,36
@@
$b_{t,\hat{g}}=$ $(1-\rho_{t})b_{t-1,\hat{g}}+\rho_{t}\Big{[}\gamma+0.5\left(\sigma^{2}_{t,g}+(\mu_{t,g}-\hat{g}_{t})^{2}\right)+0.5K_{g}\left(\sigma^{2}_{t,g}+(\mu_{t,g}-\mu_{t-1,g})^{2}\right)\Big{]},$ (36)
@@ 3.2.1 Constant VSGD and Adam
For clarity, we divide the aforementioned equation into two components: Eq. 45 and Eq. 46. The first component aligns with Eq. 39, as both represent weighted sums of an estimated second momentum and $\hat{g}_{t}^{2}$. The distinction arises in the second component, where Constant VSGD introduces two additional factors into the weighted sum: $\mu_{t-1,g}\hat{g}_{t}$ and $\tfrac{1}{K_{g}+1}\tfrac{b_{t-1,\hat{g}}}{a_{t-1}}$. Specifically, $\mu_{t-1,g}\hat{g}_{t}$ applies a penalty on Eq.45 when $\mu_{t-1,g}$ and $\hat{g}_{t}$ have opposing signs. The factor $\tfrac{1}{K_{g}+1}\frac{b_{t-1,\hat{g}}}{a_{t-1}}$ represents the $\tfrac{1}{K_{g}+1}$ fraction of observation noise learned from the data, while a higher observation noise implies greater uncertainty in $g_{t}$, leading to a higher expected value $\mathbb{E}\left[g_{t}^{2}\right]$.
In summary, Constant VSGD described in Algorithm 2 can be regarded as a variant of Adam. The first momentum update in both algorithms is equivalent when $K_{g}$ is set to $\tfrac{\beta_{1}}{1-\beta_{1}}$. However, the second momentum update in Constant VSGD incorporates an additional data-informed component that is dynamically adjusted based on the observation noise learned from the data.
@@ 3.2.2 Constant VSGD and SGD with momentum (SGDM)
$v_{t}$ $=\lambda v_{t-1}+\eta\hat{g}_{t}=(\lambda+\eta)\left[\frac{\lambda}{\lambda+\eta}v_{t-1}+\frac{\eta}{\lambda+\eta}\hat{g}_{t}\right].$ (49)
That is, the increment $v_{t}$ in SGDM is a weighted average of $v_{t-1}$ and $\hat{g}_{t}$, the counterpart of $v_{t}$ in Constant VSGD is $\mu_{t,g}$ updated by Eq. 33. Therefore, by choosing $K_{g}$ such that $K_{g}=\frac{\gamma}{\eta}$, the update equations 49 and 33 become identical except for a constant scaling factor.
@@ 3.2.3 Constant VSGD and AMSGrad
Similarly to AMSGrad, Constant VSGD incorporates a form of long-term memory regarding past gradient information in its second momentum update equations 45 and 46. However, this is not achieved by computing the maximum of historical values. Notably, the factor $\frac{1}{K_{g}+1}\frac{b_{t-1,\hat{g}}}{a_{t-1}}$ in Eq. 46 represents the expected observation noise, estimated using data up to the $(t-1)$-th iteration. In this expression, while $K_{g}$ and $a_{t-1}$ remain constant for $t\geq 2$, the variable component is $b_{t-1,\hat{g}}$, which is determined by Eq. 36, for clarity we restate it as:
@@
$s=\gamma+0.5\left(\sigma^{2}_{t,g}+(\mu_{t,g}-\hat{g}_{t})^{2}\right)+0.5K_{g}\left(\sigma^{2}_{t,g}+(\mu_{t,g}-\mu_{t-1,g})^{2}\right).$ (53)
@@ 4 Experiments
In this Section, we evaluate the performance of the VSGD compared to other optimizers on high-dimensional optimization problems  [^1] Code is available at github.com/generativeai-tue/vsgd. The task we focus on in our work is training DNNs to solve classification tasks.
4.1 Setup
Data
We used three benchmark datasets: CIFAR100 (Krizhevsky et al., 2009), TinyImagenet-200 (Deng et al., 2009a) [^2] We use Tiny Imagenet(Stanford CS231N) provided on image-net.org, and Imagenet-1k (Deng et al., 2009b). The CIFAR100 dataset contains 60000 small ($32\times 32$) RGB images labeled into 100 different classes, 50000 images are used for training, and 10000 are left for testing. In the case of TinyImagenet-200, the models are trained on 100000 images from 200 different classes and tested on 10000 images. For a large-scale experiment, we use the Imagenet-1k dataset which contains 1,281,167 images from 1000 classes.
Architectures
Hyperparameters
4.2 Results